A voltaic cell consists of an Ni/Ni2+ half-cell and a Co/Co2+ half-cell. Calculate Ecell when conc. of Ni2+ = 0.142 M and of Co2+ = 0.468 M. The reduction potential for Ni2+ is -0.247 V and for Co2+ is -0.277 V.

Answer: 0.01467

A voltaic cell consists of an Al/Al3+ half-cell and a Cd/Cd2+ half-cell. Calculate {Al3+} when {Cd2+} = 0.401 M and Ecell = 1.299 V.Use reduction potential values of Al3+ = -1.66 V and for Cd2+ = -0.40 V.

Answer: 0.003

Please help me out in these 2 questions, can you go step by step on how to solve it. Thank you!
Answer #1

Co(s) -------------> Co^2+ (aq) + 2e^-            E0 = 0.277v

Ni^2+ (aq) + 2e^- -------> Ni(s)                        E0 = -0.247v

---------------------------------------------------------------------------

Co(s) + Ni^2+ (aq) ------> Co^2+ (aq) + Ni(s)     E0cell = 0.03v

n = 2

Ecell    = E0cell - 0.0592/n logQ

             = 0.03 - 0.0592/2 log[Co^2+]/[Ni^2+]

             = 0.03 - 0.0296 log0.468/0.142

             = 0.03 - 0.0296*0.5179

              = 0.01467v >>>>answer

part-B

2Al(s) -------> 2Al^3+ (aq) + 6e^-              E0 = 1.66v

3Cd^2+ (aq) + 6e^- -----> 3Cd(s)             E0 = -0.40v

------------------------------------------------------------------------

2Al(s) + 3Cd^2+ (aq) -----> 2Al^3+ (aq) + 3Cd(s)    E0cell = 1.26v

n = 6

Ecell    = E0cell - 0.0592/n logQ

1.299         = 1.26 -0.0592/6 log[Al^3+]^2/[Cd^2+]^3

1.299        = 1.26 - 0.00986log[Al^3+]^2/(0.401)^3

1.299-1.26   = - 0.00986log[Al^3+]^2/(0.401)^3

0.039           = - 0.00986log[Al^3+]^2/(0.401)^3

log[Al^3+]^2/(0.401)^3    = 0.039/-0.00986

log[Al^3+]^2/(0.401)^3     = -3.9554

[Al^3+]^2/(0.401)^3         = 0.00011

[Al^3+]^2                      = 0.00011*(0.401)^3

[Al^3+]^2                         = 0.00000709

[Al^3+]                            = 0.003M
Answer #2

Problem 2: Calculating [Al³⁺] in a Voltaic Cell

Given:

  • Half-cells: Al/Al³⁺ (anode) and Cd/Cd²⁺ (cathode).

  • Concentrations: [Cd²⁺] = 0.401 M, [Al³⁺] = ? (to find).

  • Cell potentialEcell=1.299V.

  • Standard reduction potentials:

    • EAl³⁺/Al=1.66V.

    • ECd²⁺/Cd=0.40V.



Step 1: Determine the Standard Cell Potential (Ecell)

The cell reaction involves:

  • Oxidation (anode)AlAl³⁺+3e (Eanode=1.66V).

  • Reduction (cathode)Cd²⁺+2eCd (Ecathode=0.40V).

The standard cell potential is:

Ecell=EcathodeEanode=0.40V(1.66V)=1.26V.


Step 2: Write the Nernst Equation

The Nernst equation for the cell is:

Ecell=Ecell0.0591nlogQ,

where:

  • n=6 (total electrons transferred; balance the half-reactions: 2Al+3Cd²⁺2Al³⁺+3Cd).

  • Q=[Al³⁺]2[Cd²⁺]3.

Substitute known values:

1.299=1.260.05916log([Al³⁺]2(0.401)3).


Step 3: Solve for [Al³⁺]

  1. Rearrange the Nernst equation:

    1.2991.26=0.00985log([Al³⁺]20.0645).

  2. Simplify:

    0.039=0.00985log([Al³⁺]20.0645).

  3. Divide both sides by 0.00985:

    log([Al³⁺]20.0645)=3.96.

  4. Take the antilog:

    [Al³⁺]20.0645=103.960.000110.

  5. Solve for [Al³⁺]2:

    [Al³⁺]2=0.000110×0.06457.095×106.

  6. Take the square root:

    [Al³⁺]=7.095×1060.00266M.


Final Answer:

The concentration of [Al³⁺] is 0.003 M (rounded to 3 decimal places).


answered by: anonymous